OpenStudy (anonymous):

? What is the simplified form of 15 times x to the fifth power over 9 times y to the seventh power • 45 times y–cubed over 5 times x squared ? 15 x–cubed over y to the fourth power 15 x to the fourth power over y–cubed y to the fourth power over 15 x–cubed y–cubed over 15 x to the fourth power

OpenStudy (aaronq):

\[\frac{ (15x^5)(45y^3) }{ (9y^7)(5x^2) } = \frac{ 15x^3 }{ y^4 }\]